A pair of dice is rolled.

A pair of dice is rolled. what is the probability of getting a sum less than 9?
Answer #1


6 7 8 9 10 11 12

5 6 7 8 9 10 11

4 5 6 7 8 9 10 

3 4 5 6 7 8 9 

2 3 4 5 6 7 8

1 2 3 4 5 6 7

ANSWER: 26/36 = 13/18

answered by: Evanso mui
Answer #2

SOLUTION :


Pair of dice rolled.


P( sum < 9) = 1 - P(sum ≥ 9)


Total space = 6 * 6 = 36 events


Events,  (sum ≥ 9 ) are :


(3, 6), (6, 3), (4, 5), (5, 4), (4, 6), (6, 4), (5, 5), (5, 6), (6, 5), (6,6) = 10 events


So, P( sum ≥ 9) = 10/36 = 5/18


Hence, P (sum ≤ 9) = 1 - 5/18 = 13/18 (ANSWER).

answered by: Tulsiram Garg

> Last line be read as : Hence, P(sum < 9) = 1 - 5/18 = 13/18 . Thanks.

Tulsiram Garg Tue, Nov 16, 2021 1:46 AM